a rectangle is 1 m narrower than its long, and one of its diagonals is 2 m longer than the width of the rectangle. what is the length of the diagonal.
Answer by rachitaagarwal93(1)

let length be xm then width =(x-1)m and one of diagonal =(x-1+2)m =(x+1)m
now by pythogorous theorum digonal^2=length^2+width^2
substuting the values (x+1)^2=x^2+(x-1)^2 or x^2-4x=0 orx(x-4)=0 it gives x=4m
therefore length of rectangle=4m so width is=4-1=3m
and hence diagonal=3+2=5m